# Why Is Spatial Reasoning Hard for VLMs? An Attention Mechanism Perspective on Focus Areas

### v1

@@ Abstract
Figure 1: The framework of AdaptVis. We adaptively intervene in the temperature of the attention logits of the image tokens. 1). For generations with high confidence, we trust the attention pattern and sharpen the attention distribution. 2). For generations with low confidence, we smoothen the attention distribution to broaden the context window for better concentration on the correct objects.
@@ 1 Introduction
Despite rapid advancements in Large Vision-Language Models (VLMs), a significant deficiency persists, i.e., their struggle with vision-centric abilities (Gao et al., 2023; Kamath et al., 2023; Tong et al., 2024a; Chen et al., 2024a). This limitation is particularly notable in spatial reasoning given the simplicity of the task. Spatial reasoning involves inferring basic relationships between just two objects, such as “left”, “right”, “above”, “below”, “behind”, or “front”, as shown in Figure 1. For example, given the image with a book “behind” the candle in Figure 1, VLMs describe the book as being “left” of the candle. This error is not an isolated incident but a frequent recurring pattern that highlights a fundamental bottleneck in how VLMs process visual-centric information.
@@
While a straightforward solution might seem to be simply increasing attention to visual tokens, our deeper analysis reveals that the challenge lies not just in the quantity but in the geometric distribution of visual attention. By examining attention patterns across model layers, we observe a pattern in VLMs’ attention behavior: their initial attention distribution often reflects learned priors that may or may not align with the actual geometric distribution over the image. This led us to a key insight: rather than accepting this initial distribution, we can dynamically intervene based on the model’s self-belief, which is measured using its confidence score. As shown in Figure 1, when the model exhibits high confidence in its spatial reasoning (as measured by generation probability), we sharpen its attention distribution to strengthen focus on its current beliefs. Conversely, when confidence is low, we smooth the attention distribution to encourage exploration of alternative spatial relationships.
@@ 2 Preliminary on VLMs
Large VLMs like LLaVA (Liu et al., 2024a) consist of three components: a visual encoder like CLIP (Radford et al., 2021), a pretrained language model and a projector to connect these two parts. The visual encoder functions as a perception tool to “see” the image, while the image information is processed through a projector to be mapped into the token space. The LLM part is often based on the transformer architecture, consisting of $L$ layers stacked together. Each layer consists of two major components: a Multi-Head Attention (MHA) module, followed by a feed-forward network. For each layer $l$, given the input ${\bm{X}}\in\mathbb{R}^{n\times d}$ (where $n$ is the number of tokens and $d$ is the embedding dimension), the MHA module performs the self-attention function in each head $N_{h}$, and the output is a concatenation of all heads’ outputs: ${\text{MHA}}^{(l)}({\bm{X}})=\text{Concat}\left({\bm{N}}_{h}^{(l,1)},\dots,{\bm{N}}_{h}^{(l,H)}\right){\bm{W}}_{o},$ where $H$ is the number of heads, ${\bm{N}}_{h}$ is the output of head $N_{h}$:
${\bm{N}}_{h}^{(l,h)}=\text{Softmax}({\bm{A}}^{(l,h)}){\bm{V}}=\text{Softmax}\left(\frac{{\bm{Q}}{\bm{K}}^{\top}}{\sqrt{d_{h}}}+{\bm{M}}\right){\bm{V}}$ (1)
@@ 3 Text-Vision Attention Interactions
3.1 VLMs allocate sparse attention to the image
@@
3.2 Is the answer more accurate if the model sees the image more? Not really
@@ 4 Visual Attention Distribution
4.1 The model automatically focuses on the relevant entity when correctly answering questions
@@ 4.2 ScalingVis: Temperature scaling to image attention distribution
5 Adaptively intervene the attention distribution by model’s confidence
@@
5.1 Self-confidence reflects whether we can trust the model’s image attention distribution
@@ 5.2 AdaptVis
Recall from Section 5.1 that we observe two distinct patterns in the model’s factuality behavior: (1) synthetic data presents more unfamiliar cases than the real data, and (2) VLMs could express uncertainty through confidence scores. These insights motivate using confidence scores as a metric for adaptive intervention in the model’s internal states. Our intuition is straightforward: when confidence is low, suggesting that the attention pattern may be unreliable, we smooth the attention distribution. This encourages the model to explore a broader range of image regions, increasing the likelihood of focusing on the correct patches. Conversely, when confidence is high and attention is dispersed across the image, we sharpen the distribution to concentrate on key objects more effectively. Specifically, we apply the targeted intervention to the attention of the last input token (at the $n$-th position) directed toward the image tokens as shown in Equation 3a3b. Overall, we use (1) a large $\alpha>1$ when the Confidence $\mathcal{C}$ is large. This sharpens the attention distribution and the relevant objects are paid more attention to. (2) a small $\alpha<1$ when the confidence $\mathcal{C}$ is small. This mitigates the model’s excessive concentration on certain image tokens and makes the overall attention distribution smoother across the image.
@@
We emply a same setting with ScalingVis as shown in 4.2. For hyperparameter choice and robustness, we show in Appendix E.3.
@@ Results
Our main results are presented in Table 1 and Table 2. By controlling the distribution of attention weights, we observe a  significant improvement in spatial reasoning ability, with gains of up to 50 absolute points. In most cases, AdaptVis achieves the best performance, particularly for synthetic datasets like Cont_A and Cont_B, as shown in Table 1, where it significantly outperforms the generalized method ScalingVis. These findings suggest that model performance varies considerably with the label distribution of the dataset, and smoothing the distribution (by applying a coefficient smaller than 1) enhances performance. For the real-image datasets like COCO and VG, the adaptive method performs slightly better than the generalized approach, indicating the model’s robustness across different label distributions. Example cases are shown in Figure 9 and Figure 10, illustrating the smooth effect and focus effect ($\alpha<1$ and $\alpha>1$), respectively. It is important to note that the LLaVA-series models are trained on the COCO dataset, which makes them highly confident and familiar with COCO and VG image types. Hence trusting the model’s self-belief and sharpening image attention improves performance. Notably, for datasets containing more unfamiliar images, the adaptive setting proves to be significantly more effective.
@@ 6 Related Work
The first line of related work focuses on the attention patterns in LMs. Some studies on attention patterns in LLMs reveal biased attention across context windows, such as ineffective use of the middle context (Liu et al., 2024b) and initial token attention sinks (Xiao et al., 2023). While some approaches use fine-tuning to overcome these biases (An et al., 2024), training-free methods like input-adaptive calibration (Yu et al., 2024b) and position-specific interventions (Yu et al., 2024a) offer efficient alternatives. PASTA (Zhang et al., 2023), a closely related method, emphasizes attention on selected segments for specific heads; we extend this to VLMs without manual segment specification or multiple validation runs. Our work is also related with failure analysis in VLMs, VLMs have been shown to hallucinate more in multi-object recognition tasks and rely on spurious correlations (Chen et al., 2024c), with systematic visual limitations highlighted from a CLIP perspective (Tong et al., 2024b). Our work also connects to the decoding strategies for reducing hallucinations decoding strategies to mitigate hallucinations include contrastive decoding focusing on image regions (Leng et al., 2024), preference tuning through data augmentation (Wang et al., 2024), and methods leveraging contrastive layers for enhanced knowledge extraction (Chuang et al., 2023), as well as activation-based optimal answer identification (Chen et al., 2024b).
@@ 7 Conclusion and Future Work
Our research uncovers the inner working mechanism of VLMs during spatial reasoning, which is a critical limitation in VLMs and constrains their practical utility when requiring geometric understanding of visual scenes. We identify critical insights through an in-depth study of attention behaviors across layers: 1) VLMs allocate surprisingly insufficient attention to image tokens; 2) the location of attention on image tokens is more crucial than quantity; and 3) generation confidence serves as a reliable indicator of its familiarity with the image and the correctness of its attention pattern. Based on these findings, we propose AdaptVis, a novel decoding method that dynamically adjusts attention distribution, significantly improving spatial reasoning performance. Future research could focus on further exploring the mechanism of VLMs on complicated geometric structure understanding, such as long-horizon spatial reasoning, and investigate other spatial reasoning bottleneck.

### v2

@@ Abstract
Figure 1: The framework of AdaptVis. We adaptively intervene in the temperature of the attention logits of the image tokens. Top: For generations with low confidence, we smoothen the attention distribution to broaden the context window for better concentration on the correct objects. Bottom: For generations with high confidence, we trust the attention pattern and sharpen the attention distribution.
@@ 1 Introduction
Despite rapid advancements in Large Vision-Language Models (VLMs), a significant deficiency persists, i.e., their struggle with vision-centric abilities (Gao et al., 2023; Kamath et al., 2023; Tong et al., 2024a; Chen et al., 2024a). This limitation is particularly notable in spatial reasoning given the simplicity of the task. Spatial reasoning involves inferring basic relationships between just two objects, such as “left”, “right”, “above”, “below”, “behind”, or “front”, as shown in Figure 1. For example, given the image with a book “behind” the candle in Figure 1, VLMs describe the book as being “left” of the candle. This error is not an isolated incident but a frequent, recurring pattern, highlighting a fundamental bottleneck in how VLMs process visual-centric information.
@@
While a straightforward solution might seem to be simply increasing attention to visual tokens, our deeper analysis reveals that the challenge lies not just in the quantity but in the geometric distribution of visual attention. By examining attention patterns across model layers, we observe a pattern in VLMs’ attention behavior: their initial attention distribution often reflects learned priors that may or may not align with the actual geometric distribution over the image. This led us to a key insight: rather than accepting this initial distribution, we can dynamically intervene based on the model’s self-belief, which is measured using its confidence score. As shown in Figure 1, when the model exhibits high confidence in its spatial reasoning (as measured by generation probability), we sharpen its attention distribution to strengthen the focus on its current beliefs. Conversely, when confidence is low, we smooth the attention distribution to encourage the exploration of alternative spatial relationships.
@@ 2 Preliminary on VLMs
Large VLMs like LLaVA (Liu et al., 2024a) consist of three components: a visual encoder like CLIP (Radford et al., 2021), a pretrained language model, and a projector to connect them. The visual encoder functions as a perception tool to “see” the image, while the image information is processed through a projector to be mapped into the token space. The LLM is often based on the transformer architecture, consisting of $L$ layers stacked together. Each layer consists of two major components: a Multi-Head Attention (MHA) module, followed by a feed-forward network. For each layer $l$, given the input ${\bm{X}}\in\mathbb{R}^{n\times d}$ (where $n$ is the number of tokens and $d$ is the embedding dimension), MHA performs self-attention in each head $N_{h}$. The output is a concatenation of all heads’ outputs: ${\text{MHA}}^{(l)}({\bm{X}})=\text{Concat}\left({\bm{N}}_{h}^{(l,1)},\dots,{\bm{N}}_{h}^{(l,H)}\right){\bm{W}}_{o}.$ Here $H$ is the number of heads; ${\bm{N}}_{h}$ is the output of head $N_{h}$:
${\bm{N}}_{h}^{(l,h)}=\text{Softmax}({\bm{A}}^{(l,h)}){\bm{V}}=\text{Softmax}\left(\frac{{\bm{Q}}{\bm{K}}^{\top}}{\sqrt{d_{h}}}+{\bm{M}}\right){\bm{V}},$ (1)
@@ 3 Text-Vision Attention Interactions
3.1 VLMs allocate sparse attention to the image.
@@
3.2 Is the answer more accurate if the model sees the image more? Not really.
@@ 4 Visual Attention Distribution
4.1 The model automatically focuses on the relevant entity when correctly answering questions.
@@ 4.2 ScalingVis: Temperature scaling to image attention distribution
5 Adaptively Intervening the Attention Distribution by Model Confidence
@@
5.1 Self-confidence reflects whether we can trust the model’s image attention distribution.
@@ 5.2 AdaptVis
Recall from Section 5.1 that we observe two distinct patterns in the model’s factuality behavior: (1) synthetic data presents more unfamiliar cases than the real data, and (2) VLMs could express uncertainty through confidence scores. These insights motivate using confidence scores as a metric for adaptive intervention in the model’s internal states. Our intuition is straightforward: when confidence is low, suggesting that the attention pattern may be unreliable, we smooth the attention distribution. This encourages the model to explore a broader range of image regions, increasing the likelihood of focusing on the correct patches. Conversely, when confidence is high and attention is dispersed across the image, we sharpen the distribution to concentrate on key objects more effectively. Specifically, we apply the targeted intervention to the attention of the last input token (at the $n$-th position) directed toward the image tokens as shown in Equation 3a3b. Overall, we use a large $\alpha>1$ when the Confidence $\mathcal{C}$ is large, which sharpens the attention distribution, and the relevant objects are paid more attention to; we use a small $\alpha<1$ when the confidence $\mathcal{C}$ is small, which mitigates the model’s excessive concentration on certain image tokens and makes the overall attention distribution smoother across the image.
@@
We employ the same setting with ScalingVis as shown in 4.2. For hyperparameter choice and robustness, we show in Appendix E.3.
@@ Results
Our main results are presented in Table 1 and Table 2. By controlling the distribution of attention weights, we observe a  significant improvement in spatial reasoning ability, with gains of up to 50 absolute points. In most cases, AdaptVis achieves the best performance, particularly for synthetic datasets like Cont_A and Cont_B, as shown in Table 1, where it significantly outperforms the generalized method ScalingVis. These findings suggest that model performance varies considerably with the label distribution of the dataset, and smoothing the distribution (by applying a coefficient smaller than 1) enhances performance. For real-image datasets like COCO and VG, the adaptive method performs slightly better than the generalized approach, indicating the model’s robustness across different label distributions. Example cases are shown in Figure 9 and Figure 10, illustrating the smooth effect and focus effect ($\alpha<1$ and $\alpha>1$), respectively. It is important to note that the LLaVA-series models are trained on the COCO dataset, which makes them highly confident and familiar with COCO and VG image types. Hence trusting the model’s self-belief and sharpening image attention improves performance. Notably, for datasets containing more unfamiliar images, the adaptive setting proves to be significantly more effective.
@@ 6 Related Work
The first line of related work focuses on the attention patterns in LMs. Some studies on attention patterns in LLMs reveal biased attention across context windows, such as ineffective use of the middle context (Liu et al., 2024b) and initial token attention sinks (Xiao et al., 2023). While some approaches use fine-tuning to overcome these biases (An et al., 2024), training-free methods like input-adaptive calibration (Yu et al., 2024b) and position-specific interventions (Yu et al., 2024a) offer efficient alternatives. PASTA (Zhang et al., 2023), a closely related method, emphasizes attention on selected segments for specific heads; we extend this to VLMs without manual segment specification or multiple validation runs. Our work is also related to failure analysis in VLMs, VLMs have been shown to hallucinate more in multi-object recognition tasks and rely on spurious correlations (Chen et al., 2024c), with systematic visual limitations highlighted from a CLIP perspective (Tong et al., 2024b). Our work also connects to the decoding strategies for reducing hallucinations decoding strategies to mitigate hallucinations include contrastive decoding focusing on image regions (Leng et al., 2024), preference tuning through data augmentation (Wang et al., 2024), and methods leveraging contrastive layers for enhanced knowledge extraction (Chuang et al., 2023), as well as activation-based optimal answer identification (Chen et al., 2024b).
@@ 7 Conclusion and Future Work
Our research uncovers the inner working mechanism of VLMs during spatial reasoning, which is a critical limitation in VLMs and constrains their practical utility when requiring geometric understanding of visual scenes. We identify critical insights through an in-depth study of attention behaviors across layers: 1) VLMs allocate surprisingly insufficient attention to image tokens; 2) the location of attention on image tokens is more crucial than quantity; and 3) generation confidence serves as a reliable indicator of its familiarity with the image and the correctness of its attention pattern. Based on these findings, we propose AdaptVis, a novel decoding method that dynamically adjusts attention distribution, significantly improving spatial reasoning performance. Future research could focus on further exploring the mechanism of VLMs on complicated geometric structure understanding, such as long-horizon spatial reasoning, and investigate other spatial reasoning bottlenecks.